# Magnetic Field Amplification and Particle Acceleration in Weakly Magnetized Trans-relativistic Electron-ion Shocks

### v1

@@ Appendix A Unmagnetized Trans-Relativistic Shock
With the above caveats in mind, we present here, for completeness, the unmagnetized case. Figure 10 shows a snapshot of an unmagnetized $\sigma=0$ shock taken at $\omega_{\mathrm{pi}}t=5100$. The simulation parameters are the same as described in the main text (Section 2). In panels (a,b), we see magnetized plasma cavities that are noticeably larger than the ion skin depth scale. The role of plasma cavities has been extensively discussed in recent works, focusing on the ultra-relativistic regime (Peterson2022; Groselj2024; Demidov2025). Here, we find that plasma cavities can also be generated in trans-relativistic unmagnetized shocks. As found in the ultra-relativistic regime, these cavities may enable efficient field generation and particle acceleration. The behavior seen in Figure 10 may be representative of Weibel-dominated shocks at magnetizations smaller than the range covered in the main text, e.g., $\sigma\sim 10^{-9}$, as appropriate for the interstellar medium.
@@ Appendix B Saturation Level of the Bell Instability
Two scalings have been proposed in the literature regarding the saturation of Bell modes. The first one pertains to the low-current regime and dictates that the magnetic field energy saturates at a fixed fraction of the cosmic ray momentum flux. The idea originates from the pioneering work by Bell2004. Zacharegkas2024 have recently formulated the saturation mechanism in a more precise manner, and validated their predictions by means of hybrid-kinetic simulations. They propose
@@
Recent works showed that the nature of Bell modes in the high-current regime $\eta>\eta_{\mathrm{crit}}$ is different (Weidl2019; Lichko2025). Here, the magnetic energy density of Bell modes saturates at a fixed fraction of the ambient field energy. This is common among gyro-resonant plasma instabilities (Stix1992). Lichko2025 found that $\delta B^{2}/B_{0}^{2}\sim 25$ for the high-current regime of the Bell instability.
@@
We perform 1D periodic box PIC simulations to clarify the saturation level in the parameter regime of interest, which can be seen as a trans-relativistic generalization of previous studies. We set up the simulation in the upstream frame with a background ion density of $n_{0}$. The ambient field $\bm{B}_{0}$ is in the $x$-direction, along the simulation box. Definitions of the background plasma quantities, $\omega_{\mathrm{p}s},\Omega_{s},d_{s}$, and $\sigma$ are the same as in the main text. The cosmic ray component has a proper density of $n_{\mathrm{CR}}$, drifting in the positive $x$-direction with velocity $v_{\mathrm{d}}$. Note that the simulation frame number density is $\gamma_{\mathrm{d}}n_{\mathrm{CR}}$, and the cosmic ray current (normalized to $en_{0}c$) is $\eta=\beta_{\mathrm{d}}\gamma_{\mathrm{d}}n_{\mathrm{CR}}/n_{0}$. For charge and current neutrality, the background electrons are initialized with a simulation frame density of $n_{0}+\gamma_{\mathrm{d}}n_{\mathrm{CR}}$ and drift velocity of $v_{\mathrm{d}}\gamma_{\mathrm{d}}n_{\mathrm{d}}/(n_{0}+\gamma_{\mathrm{d}}n_{\mathrm{CR}})$. We fix the drift velocity $\beta_{\mathrm{d}}=0.8$, and investigate the dependence on $\eta,\gamma_{\mathrm{iso}}$ and $\sigma$. The simulation parameters are $\Delta x/d_{\mathrm{e}}=0.1,c\Delta t/\Delta x=0.99$, and $L_{x}=153.6d_{\mathrm{i}}$. We use $N_{\mathrm{ppc}}=1024$ particles per cell, with a quartic spline shape function, for each of the three components, which are background ions, cosmic ray ions, and electrons. We use the open source code SMILEI (Derouillat2018) in this Appendix.
@@
Let us start with the low-current regime. Figure 11 shows the results for fixed magnetization $\sigma=10^{-3}$ and current $\eta=0.02$, resulting in $\eta/\eta_{\mathrm{crit}}=0.32$. We vary the cosmic ray energy $\gamma_{\mathrm{iso}}=[2.5,80]$. The time evolution of the spatially averaged magnetic field energy $\delta B^{2}=B_{y}^{2}+B_{z}^{2}$ is shown with two different normalizations. Panel (a) uses $8\pi\beta_{\mathrm{d}}T^{01}_{\mathrm{CR}}$ normalization for the magnetic field energy and $\omega_{\mathrm{pi}}$ for the time. The growth rate is consistent with the theoretical prediction $\Gamma/\omega_{\mathrm{pi}}=\eta/2$ (Bell2004; Amato2009). At the peak $\omega_{\mathrm{pi}}t\sim 1500$, we measure the saturation level of $\delta B^{2}/8\pi\beta_{\mathrm{d}}T^{01}_{\mathrm{CR}}$ in the range of $[0.1,0.4]$, which is consistent with the value $\sim 1/4$ by Zacharegkas2024. However, we see a systematic $\gamma_{\mathrm{iso}}$-dependence of $\delta B^{2}/8\pi\beta_{\mathrm{d}}T_{\mathrm{CR}}^{01}\propto\gamma_{\mathrm{iso}}^{-0.2}$, which was not discussed in their work. This is most likely due to relativistic effects absent in hybrid simulations. Panel (b) uses $B^{2}_{0}$ and $\Omega_{\mathrm{i}}$ for normalization of magnetic energy and time, respectively. We can see that the magnetic field amplification factor increases with larger cosmic ray energy, with the scaling $\delta B^{2}/B_{0}^{2}\propto\gamma_{\mathrm{iso}}^{0.8}$. Again, this is a slightly weaker scaling than $\delta B^{2}/B_{0}^{2}\propto\gamma_{\mathrm{iso}}$ as proposed by Zacharegkas2024. Field amplification proportional to cosmic-ray energy presents favorable conditions for the long-term acceleration of particles at low-current Bell-dominated shocks.
@@
Figure 12 shows the results for the high-current regime. The magnetization and the current are $\sigma=10^{-4}$ and $\eta=0.1$, resulting in $\eta/\eta_{\mathrm{crit}}=5$. The growth rate is consistent with the high-current Bell theoretical prediction $\Gamma=\Omega_{\mathrm{i}}$ (Weidl2019). In panel (b), we can see a very consistent $\delta B^{2}/B_{0}^{2}\simeq 100$. This is comparable with the results by Lichko2025 with a factor 4 difference. When normalized by the cosmic ray momentum flux (panel (a)), it is apparent that high-current Bell modes cannot efficiently tap into the cosmic ray energy, especially at high $\gamma_{\mathrm{iso}}$.
@@
Figure 12 and the results by Lichko2025 elucidate the saturation level of the high-current Bell instability for a fixed $\sigma$. To fully understand the transition from a Bell-dominated shock to a Weibel-dominated shock, as we saw in the main text (Subsection 3.1), we also need to clarify the dependence on the magnetization. Figure 13 shows the result for fixed $\gamma_{\mathrm{iso}}=10$ and $\eta=0.1$, while varying $\sigma$. Panel (a), in which the fluctuating field energy is normalized to the cosmic ray momentum flux, shows a clear decline in the saturation level at lower magnetizations. In panel (b), we see that the growth rate is consistently $\Gamma=\Omega_{\mathrm{i}}$. For the saturation level, it was expected that the $\sigma=10^{-3}$ case, in dark teal, would perform differently, since $\eta/\eta_{\mathrm{crit}}$ is only slightly above unity, at $1.6$. At magnetizations below $\sigma=10^{-3}$, which are significantly supercritical $(\eta/\eta_{\mathrm{crit}}\gg 1)$, a similar $\delta B^{2}/B_{0}^{2}\sim 100$ magnetic field saturation level as in Figure 12 is found. Although we see some variation in the saturation level, we find that there is no systematic $\sigma$-dependence when the instability operates in the high-current regime.
@@ Appendix C Self-Regulation of Cosmic Ray Current and Energy Partition
In the main text (Section 3), we saw that the strength of the cosmic ray current has a significant impact on shock structure and particle acceleration. Here, we discuss the different types of shock configurations that can be envisioned in the trans-relativistic regime, as a result of nonlinear feedback between accelerated particles and their self-generated fields.
@@
[…]
$\frac{r_{\mathrm{g,e}}}{d_{\mathrm{i}}}=7\left(\frac{\sigma}{10^{-3}}\right)^{-1/2}\left(\frac{\delta B/B_{0}}{10}\right)^{-1}\left(\frac{u_{\mathrm{ref}}/c}{7}\right)\left(\frac{[m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}}{10}\right)^{-1/2}.$ (C1)
[…]
We have made the following assumptions. We take $\delta B/B_{0}\sim 10$ as a universal value for the high-current Bell regime appropriate for state IIa (Appendix B). We assume elastic reflection at the shock front, which leads to a typical four-velocity after reflection of $u_{\mathrm{ref}}\sim\Gamma_{\mathrm{sh}}^{2}V_{\mathrm{sh}}\sim 7c$. The mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}$ is $\sim 1836$ for protons for realistic parameters, and 100 in our PIC simulations. However, accounting for electron heating, the effective mass ratio becomes $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 10$ (see Subsection 3.3). If this Larmor radius is smaller than the typical half wavelength of high-current Bell modes, $\lambda/2d_{\mathrm{i}}\sim 10$, which we assume to be independent of the initial $m_{\mathrm{i}}/m_{\mathrm{e}}$ (Figure 3(c)), electron reflection is suppressed. Note that electrons are always affected more dramatically than ions, due to their smaller Larmor radius. If this happens, the shock downstream becomes negatively charged, and the electrostatic field pulls back some of the returning ions towards the downstream. This reduces the cosmic ray current, resulting in the transition to an ion-dominated low-current cosmic ray state, that is, state I. This is consistent with the time evolution of the $\sigma=10^{-3}$ case in the main text. The dark teal curve in Figure 5 reaches a high-current state at $\omega_{\mathrm{pi}}t\sim 2000$, but self-regulates to a low-current state at $\omega_{\mathrm{pi}}t\gtrsim 5000$.

### v2

@@ Appendix A Unmagnetized Trans-Relativistic Shock
With the above caveats in mind, we present here, for completeness, the unmagnetized case. Figure 10 shows a snapshot of an unmagnetized $\sigma=0$ shock taken at $\omega_{\mathrm{pi}}t=5100$. The simulation parameters are the same as described in the main text (Section II). In panels (a,b), we see magnetized plasma cavities that are noticeably larger than the ion skin depth scale. The role of plasma cavities has been extensively discussed in recent works, focusing on the ultra-relativistic regime (Peterson et al., 2022; Grošelj et al., 2024; Demidov et al., 2026). Here, we find that plasma cavities can also be generated in trans-relativistic unmagnetized shocks. As found in the ultra-relativistic regime, these cavities may enable efficient field generation and particle acceleration. The behavior seen in Figure 10 may be representative of Weibel-dominated shocks at magnetizations smaller than the range covered in the main text, e.g., $\sigma\sim 10^{-9}$, as appropriate for the interstellar medium.
@@ Appendix B Saturation Level of the Bell Instability
Two scalings have been proposed in the literature regarding the saturation of Bell modes. The first one pertains to the low-current regime and dictates that the magnetic field energy saturates at a fixed fraction of the cosmic ray momentum flux. The idea originates from the pioneering work by Bell (2004). Zacharegkas et al. (2024) have recently formulated the saturation mechanism in a more precise manner, and validated their predictions by means of hybrid-kinetic simulations. They propose
@@
Recent works showed that the nature of Bell modes in the high-current regime $\eta>\eta_{\mathrm{crit}}$ is different (Weidl et al., 2019; Lichko et al., 2025). Here, the magnetic energy density of Bell modes saturates at a fixed fraction of the ambient field energy. This is common among gyro-resonant plasma instabilities (Stix, 1992). Lichko et al. (2025) found that $\delta B^{2}/B_{0}^{2}\sim 25$ for the high-current regime of the Bell instability.
@@
We perform 1D periodic box PIC simulations to clarify the saturation level in the parameter regime of interest, which can be seen as a trans-relativistic generalization of previous studies. We set up the simulation in the upstream frame with a background ion density of $n_{0}$. The ambient field $\bm{B}_{0}$ is in the $x$-direction, along the simulation box. Definitions of the background plasma quantities, $\omega_{\mathrm{p}s},\Omega_{s},d_{s}$, and $\sigma$ are the same as in the main text. The cosmic ray component has a proper density of $n_{\mathrm{CR}}$, drifting in the positive $x$-direction with velocity $v_{\mathrm{d}}$. Note that the simulation frame number density is $\gamma_{\mathrm{d}}n_{\mathrm{CR}}$, and the cosmic ray current (normalized to $en_{0}c$) is $\eta=\beta_{\mathrm{d}}\gamma_{\mathrm{d}}n_{\mathrm{CR}}/n_{0}$. For charge and current neutrality, the background electrons are initialized with a simulation frame density of $n_{0}+\gamma_{\mathrm{d}}n_{\mathrm{CR}}$ and drift velocity of $v_{\mathrm{d}}\gamma_{\mathrm{d}}n_{\mathrm{d}}/(n_{0}+\gamma_{\mathrm{d}}n_{\mathrm{CR}})$. We fix the drift velocity $\beta_{\mathrm{d}}=0.8$, and investigate the dependence on $\eta,\gamma_{\mathrm{iso}}$ and $\sigma$. The simulation parameters are $\Delta x/d_{\mathrm{e}}=0.1,c\Delta t/\Delta x=0.99$, and $L_{x}=153.6d_{\mathrm{i}}$. We use $N_{\mathrm{ppc}}=1024$ particles per cell, with a quartic spline shape function, for each of the three components, which are background ions, cosmic ray ions, and electrons. We use the open source code SMILEI (Derouillat et al., 2018) in this Appendix.
@@
Let us start with the low-current regime. Figure 11 shows the results for fixed magnetization $\sigma=10^{-3}$ and current $\eta=0.02$, resulting in $\eta/\eta_{\mathrm{crit}}=0.32$. We vary the cosmic ray energy $\gamma_{\mathrm{iso}}=[2.5,80]$. The time evolution of the spatially averaged magnetic field energy $\delta B^{2}=B_{y}^{2}+B_{z}^{2}$ is shown with two different normalizations. Panel (a) uses $8\pi\beta_{\mathrm{d}}T^{01}_{\mathrm{CR}}$ normalization for the magnetic field energy and $\omega_{\mathrm{pi}}$ for the time. The growth rate is consistent with the theoretical prediction $\Gamma/\omega_{\mathrm{pi}}=\eta/2$ (Bell, 2004; Amato and Blasi, 2009). At the peak $\omega_{\mathrm{pi}}t\sim 1500$, we measure the saturation level of $\delta B^{2}/8\pi\beta_{\mathrm{d}}T^{01}_{\mathrm{CR}}$ in the range of $[0.1,0.4]$, which is consistent with the value $\sim 1/4$ by Zacharegkas et al. (2024). However, we see a systematic $\gamma_{\mathrm{iso}}$-dependence of $\delta B^{2}/8\pi\beta_{\mathrm{d}}T_{\mathrm{CR}}^{01}\propto\gamma_{\mathrm{iso}}^{-0.2}$, which was not discussed in their work. This is most likely due to relativistic effects absent in hybrid simulations. Panel (b) uses $B^{2}_{0}$ and $\Omega_{\mathrm{i}}$ for normalization of magnetic energy and time, respectively. We can see that the magnetic field amplification factor increases with larger cosmic ray energy, with the scaling $\delta B^{2}/B_{0}^{2}\propto\gamma_{\mathrm{iso}}^{0.8}$. Again, this is a slightly weaker scaling than $\delta B^{2}/B_{0}^{2}\propto\gamma_{\mathrm{iso}}$ as proposed by Zacharegkas et al. (2024). Field amplification proportional to cosmic-ray energy presents favorable conditions for the long-term acceleration of particles at low-current Bell-dominated shocks.
@@
Figure 12 shows the results for the high-current regime. The magnetization and the current are $\sigma=10^{-4}$ and $\eta=0.1$, resulting in $\eta/\eta_{\mathrm{crit}}=5$. The growth rate is consistent with the high-current Bell theoretical prediction $\Gamma=\Omega_{\mathrm{i}}$ (Weidl et al., 2019). In panel (b), we can see a very consistent $\delta B^{2}/B_{0}^{2}\simeq 100$. This is comparable with the results by Lichko et al. (2025) with a factor 4 difference. When normalized by the cosmic ray momentum flux (panel (a)), it is apparent that high-current Bell modes cannot efficiently tap into the cosmic ray energy, especially at high $\gamma_{\mathrm{iso}}$.
@@
Figure 12 and the results by Lichko et al. (2025) elucidate the saturation level of the high-current Bell instability for a fixed $\sigma$. To fully understand the transition from a Bell-dominated shock to a Weibel-dominated shock, as we saw in the main text (Subsection III.1), we also need to clarify the dependence on the magnetization. Figure 13 shows the result for fixed $\gamma_{\mathrm{iso}}=10$ and $\eta=0.1$, while varying $\sigma$. Panel (a), in which the fluctuating field energy is normalized to the cosmic ray momentum flux, shows a clear decline in the saturation level at lower magnetizations. In panel (b), we see that the growth rate is consistently $\Gamma=\Omega_{\mathrm{i}}$. For the saturation level, it was expected that the $\sigma=10^{-3}$ case, in dark teal, would perform differently, since $\eta/\eta_{\mathrm{crit}}$ is only slightly above unity, at $1.6$. At magnetizations below $\sigma=10^{-3}$, which are significantly supercritical $(\eta/\eta_{\mathrm{crit}}\gg 1)$, a similar $\delta B^{2}/B_{0}^{2}\sim 100$ magnetic field saturation level as in Figure 12 is found. Although we see some variation in the saturation level, we find that there is no systematic $\sigma$-dependence when the instability operates in the high-current regime.
@@ Appendix C Self-Regulation of Cosmic Ray Current and Energy Partition
In the main text (Section III), we saw that the strength of the cosmic ray current has a significant impact on shock structure and particle acceleration. Here, we discuss the different types of shock configurations that can be envisioned in the trans-relativistic regime, as a result of nonlinear feedback between accelerated particles and their self-generated fields.
@@
Now we argue that the shock configuration corresponding to state II is generally unstable, meaning that the shock will transition to either state I or III. Consider a population of shock-reflected electrons with energies comparable to the typical downstream energy, under state IIa. Their Larmor radius in the amplified field can be written as
$\frac{r_{\mathrm{g,e}}}{d_{\mathrm{i}}}=$ $7\left(\frac{\sigma}{10^{-3}}\right)^{-1/2}\left(\frac{\delta B/B_{0}}{10}\right)^{-1}$ (C1)
$\times\left(\frac{u_{\mathrm{ref}}/c}{7}\right)\left(\frac{[m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}}{10}\right)^{-1/2}.$
We have made the following assumptions. We take $\delta B/B_{0}\sim 10$ as a universal value for the high-current Bell regime appropriate for state IIa (Appendix B). We assume elastic reflection at the shock front, which leads to a typical four-velocity after reflection of $u_{\mathrm{ref}}\sim\Gamma_{\mathrm{sh}}^{2}V_{\mathrm{sh}}\sim 7c$. The mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}$ is $\sim 1836$ for protons for realistic parameters, and 100 in our PIC simulations. However, accounting for electron heating, the effective mass ratio becomes $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 10$ (see Subsection III.3). If this Larmor radius is smaller than the typical half wavelength of high-current Bell modes, $\lambda/2d_{\mathrm{i}}\sim 10$, which we assume to be independent of the initial $m_{\mathrm{i}}/m_{\mathrm{e}}$ (Figure 3(c)), electron reflection is suppressed. Note that electrons are always affected more dramatically than ions, due to their smaller Larmor radius. If this happens, the shock downstream becomes negatively charged, and the electrostatic field pulls back some of the returning ions towards the downstream. This reduces the cosmic ray current, resulting in the transition to an ion-dominated low-current cosmic ray state, that is, state I. This is consistent with the time evolution of the $\sigma=10^{-3}$ case in the main text. The dark teal curve in Figure 5 reaches a high-current state at $\omega_{\mathrm{pi}}t\sim 2000$, but self-regulates to a low-current state at $\omega_{\mathrm{pi}}t\gtrsim 5000$.